# On Inverse Problems, Parameter Estimation, and Domain Generalization

### v1

@@ 2.4 Cramér-Rao Bound
[…]
@@
$E(\theta-\hat{\theta}_{X})^{2}\geq\frac{1}{J_{X}(\theta)},$ (5)
In other words, the Fisher information of the likelihood function fundamentally bounds the precision to which we can estimate $\theta$.
@@ 2.5 Problem Formulation
Denote $Y$ as an observation, and $\hat{\theta}_{Y}$ as the estimated parameter given $m$ examples (realizations) of $Y$. We assume $\hat{X}=\mathcal{F}(Y)$ is an estimated variable that is derived from $Y$ by solving an inverse problem, in an attempt to recover an underlying source signal $X$, and $\hat{\theta}_{\hat{X}}$ as the estimated parameter given $m$ examples of $\hat{X}$. In other words, $\theta-X-Y$ is a Markov chain. That is, given $X$, $Y$ and $\theta$ are independent. Namely, $p_{\theta}(x,y)\triangleq p(x,y|\theta)=p_{\theta}(x)p(y|x)$ (i.e., the conditional distribution of $Y$ given $X$ is independent of $\theta$).
@@
Definition 2 (see (Ohayon et al., 2023b), Definition 3.1)
A degradation is invertible if $p_{X|Y}(\cdot|y)$ is a delta function for almost every $y\in\mathrm{supp}(p_{Y})$.
@@
Figure 1: Illustration of an example of the problem setting: Classification with or without image-to-image regression. We observe a degraded image $y\sim p_{Y|\theta}$ according to some conditional distribution $p_{Y|X}$, where the original image is a realization from a distribution $x\sim p_{X|\theta}$.
@@ 3 Towards mathematical theory of domain variation
The Domain-Shift Problem. In learning theory, domain shift is a variation of the data distribution between the source domain (training dataset) and target domain (the test dataset). Supervised learning models have demonstrated remarkable results in image-to-image translation tasks across numerous applications (Elad et al., 2023; Pereg, 2023b, 2025). However, it has been observed that while sufficient results were obtained for the training domain (source domain), these models collapse when tested on unseen target domains, such as under contrast and resolution variations (Panigrahi et al., 2021). One of the most common techniques to improve generalization is data augmentation, normally accomplished by applying simple spatial and intensity transforms to the training data. Another popular strategy to improve generalization is domain randomization (Tobin et al., 2017) attempting to significantly enlarge the training distribution beyond the source domain, under the hypothesis that learning from out-of-distribution examples improves generalization (Bengio et al., 2011). Most domain adaptation and transfer learning methods operate under the assumption that there exists a shared hypothesis that performs well in both domains (Ben-David et al., 2010). Here, we argue that certain implementations of this strategy may lead to degraded results, because instead of learning out-of-distribution examples (extrapolation) the model learns new examples within the distribution’s support that compete with the desired task and create ambiguity, which in turn leads to mode-collapse.
@@ 3.1 Signal Restoration over Several Domains
$p_{\theta_{i}}(\mathbf{x}_{i}|\mathbf{y})=\delta(\mathbf{x}_{i}-G_{i}(\mathbf{y})),\ G_{i}(\mathbf{y})\neq G_{j}(\mathbf{y})\ \ \forall i\neq j,\ i,j=1,...,M,$ (6)
$G_{i}(\mathbf{y})$ is a deterministic function. In other words, in different domains the forward model $\mathcal{T}:\mathbf{x}\rightarrow\mathbf{y}$ (random or deterministic) leads different source signals to the same observed measurement. In different domains the observed measurement originates in different source signals. We are training an algorithm in a supervised fashion to produce a deterministic mapping $\mathcal{F}:\mathbf{y}\rightarrow\mathbf{x}$ such that $p_{\hat{X}|Y}(\hat{\mathbf{x}}|\mathbf{y})=\delta\big{(}\mathcal{F}(\mathbf{y})-\hat{\mathbf{x}}\big{)}$. When adopting domain randomization approach  [^1] We note that domain randomization inherently assumes that all domains are equally probable. That is, $P(\theta_{i})=\frac{1}{M}\forall\ i$. And one may wonder if this is indeed the best approach?, or a data augmentation approach  [^2] Under a data augmentation strategy that generates examples from domains out of the source domain., we are optimizing a neural net’s parameters by minimizing some loss function $\mathcal{L}(\hat{\mathbf{x}},\mathbf{x}_{i})$ that has the smallest average deviation from the expected outputs. But since we have $M$ domains, i.e., $M$ valid outputs, we now have
$\arg\min_{\hat{\mathbf{x}}}\sum_{i=1}^{M}{\ell(\hat{\mathbf{x}},\mathbf{x}_{i})},$ (7)
@@
$\hat{\mathbf{x}}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{x}_{i}.$ (8)
For the $\ell_{1}$ loss, the sum of absolute deviations $\ell(\hat{\mathbf{x}},\mathbf{x}_{i})=|\hat{\mathbf{x}}-\mathbf{x}_{i}|$ the optimum is at the median of the expected outputs. This phenomenon is also referred to in the literature as “regression to the mean” (Elad et al., 2023), but normally in the context of having multiple restoration solutions in one domain due the ill-posed nature of the inversion task (see also Section 4.3.2 below). Here, we emphasize that even if the restoration function is invertible in each domain  [^3] That is, in each domain the posterior is a delta function with nonzero probability for only one reconstruction. Therefore it is possible to perfectly reconstruct the signal from the measurement. unless the sets are disjoint, domain randomization would directly lead to averaging over the possible outputs. Let us summarize the above conclusion in the following theorem,
Theorem 3 (Double Meaning Theorem)
Given M domains $p(\mathbf{x}|\theta_{i}),\ i=1,...,M$ and a degradation process $p(\mathbf{y}|{\mathbf{x}})$, where the observation sets are not disjoint, i.e., $p_{\theta_{i}}(\mathbf{x}|\mathbf{y})$ have joint supports in the $\mathbf{y}$ input domain. Assuming an inversion model that is successfully trained by domain randomization, that is, the model is trained to minimize a cost function over all given and equiprobable domains, then the inversion output is expected to be such that it yields the minimal average expected loss over the possible outputs in the different domains. Mathematically speaking,
$\hat{\mathbf{x}}=\mathcal{F}(\mathbf{y})=\arg\min_{\hat{\mathbf{x}}}\frac{1}{M}\sum_{i=1}^{M}{\ell(\hat{\mathbf{x}},\mathbf{x}_{i})}.$ (9)
Proof. See above.
@@
Remark 2. Note that when the sets $\{\mathbf{y}\sim p_{\theta_{i}}(\mathbf{y})\}_{i=1}^{M}$ are disjoint, it is possible to train a model in a randomized manner over these distributions to generalize over the general $p(\mathbf{x}|\mathbf{y})$ without yielding averaged results. This may be the case for certain implementations of segmentation over different domains (Dey et al., 2024), and for the MNIST dataset that can be modeled as a Gaussiam mixture model, under the assumption that the Gaussians are “far enough” from each other.
@@
Remark 4. A similar proof outline applies also when a model is trained with noisy or perturbed training data (which can be treated as an adversarial attack (Sofer et al., 2025)) and for other regression tasks. In those cases the model converges to the mean, which, depending on the task is either beneficial or misleading.
Simple is beautiful. In previous works, it has also been observed that synthetic data training may improve generalization (Pereg et al., 2020a; Dey et al., 2024; Pereg, 2025). In those works, the synthetic data is of simpler visual form, almost degenerated. This may be explained by the opposing approach. That is, one may claim that using synthetic data of extremely simplified forms we are significantly limiting the learned distribution, and rather than expanding it, we are contracting the example set which the model has seen and is able to recognize. And so, by accurately learning simpler representations - the general result is improved. The model learns the basic pattern that may align with many distributions, inherently learning the overlapping support shared across a greater number of possible varying distributions. An alternative theoretical explanation relying on the information theoretic method of types is presented in (Pereg, 2023a).
We now provide an illustration of the core idea in the Double Meaning Theorem (Theorem 3) in two relatively simple yet pragmatic examples.
@@ 3.2 Example 1 - Resolution Shift
Assume a source image and a degraded image in vector form, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{N\times 1}$, such that, $\{\mathbf{x},\mathbf{y}\sim p_{\theta_{i}}(\mathbf{x},\mathbf{y}):\mathbf{y}=\mathbf{H}_{i}\mathbf{x}+\mathbf{w},\ i=\{1,2\}\}$. $\mathbf{H}_{i}$ is the convolution matrix of a Gaussian filter with standard deviation $\sigma_{i}$, where $\sigma_{2}=2\sigma_{1}=2$, and $\mathbf{w}$ is additional white Gaussian noise (WGN) of level $\sigma_{n}=\sqrt{2}$. To validate our theoretical results, we compare trained one-shot supervised learning models for the task of image deblurring (Pereg, 2023b, 2025). Recall that for a Gaussian PSF, the convolution of two Gaussian with mean $\mu_{1},\mu_{2}$ and variance $\sigma^{2}_{1}$, $\sigma^{2}_{2}$ is a Gaussian with mean $\mu=\mu_{1}+\mu_{2}$ and variance $\sigma^{2}=\sigma^{2}_{1}+\sigma^{2}_{2}$. Therefore, under this model, there exist $\mathbf{x}_{1}\neq\mathbf{x}_{2}$ such that $\mathbf{y}=\mathbf{H}_{1}\mathbf{x}_{1}=\mathbf{H}_{2}\mathbf{x}_{2}$. Let us denote $\mathbf{y}_{1}=\mathbf{H}_{1}\mathbf{x}_{1},\ \mathbf{y}_{2}=\mathbf{H}_{2}\mathbf{x}_{2}$, where we have omitted the noise term for that sake of simplicity. An identical degraded input image obeys $\mathbf{y}_{1}=\mathbf{y}_{2}$. Therefore,
$\mathbf{x}_{1}*\mathbf{h}_{1}=\mathbf{x}_{2}*\mathbf{h}_{2}=\mathbf{x}_{2}*\mathbf{h}_{1}*\mathbf{h}_{1\rightarrow 2},$ (10)
where $\mathbf{h}_{1\rightarrow 2}$ is a Gaussian filter with $\sigma^{2}_{1\rightarrow 2}=\sigma_{2}^{2}-\sigma_{1}^{2}$. That is, $\mathbf{x}_{1}=\mathbf{x}_{2}*\mathbf{h}_{1\rightarrow 2}$. Consequently, according to Theorem 3, the outcome of a model that optimizes the MSE over both outcomes and thus averaging over both viable reconstructions is,
$\hat{\mathbf{x}}=\frac{1}{2}\mathbf{x}_{1}+\frac{1}{2}\mathbf{x}_{2}=\frac{1}{2}\mathbf{x}_{2}+\frac{1}{2}\mathbf{x}_{2}*\mathbf{h}_{1\rightarrow 2}=\frac{1}{2}[\mathbf{I+H}_{1\rightarrow 2}]\mathbf{x}_{2}.$ (11)
@@
To illustrate the Double Meaning theorem above, we present a simple example using a one-shot training framework previously described in (Pereg, 2023b, 2025). Figures 3-4(a)-(b) show the source image and the degraded image for $\sigma_{2}=2$, respectively. The inverted results and the residuals (absolute value) are depicted in figures 3-4(c)-(d) and figures 3-4(e)-(f), respectively. We compare the result of a model that was trained with images of both degradation types (mixed training) with a model that was trained solely for the Gaussian blur with a matching $\sigma_{2}=2$ (targeted training). As can be visually observed, in accordance with the Double Meaning Theorem, mixed training yields inferior results compared with targeted training where the training matches with the blurring kernel. As can be seen, as expected, the model cannot adhere to both required resolution levels and it “collapses to the mean”.
@@ 3.3 Example 2 - Sampling Shift
Similarly, we investigate the generalization for speckle-suppression in optical coherence tomography (Pereg, 2023b), and compare the ability of the model to generalize over images of two lateral sampling rates $f^{1}_{\mathrm{s}}=2f^{1}_{\mathrm{s}}$. We test a model that is trained with retinal images of both sampling rates, and a single matching sampling rate. Figures 5-6 show the speckled images, despeckled images obtained by applying Non-Local-Means (NLM) filtering (Cuartas-Vélez et al., 2018) used as ground truth, the despeckled images produced by the trained models, and the residuals. Clearly, mixing sampling rates during training yields degraded results. Note the increased error around the blood vessel shadow projections (darker vertical area) in Fig. 5. The blood vessel shadow is almost smeared completely in the mixed training example. We observe that, as expected, the models’ prediction is more loyal to the measurement than to the NLM ground truth. Additionally, we can observe crossed-hatched artifacts as a result of inconsistencies in training sampling rates. For these examples the differences between the outputs may often seem subtle, because the averaged terms are visually similar and their average may also seem close to the ground truth.

### v2

@@ 2.4 Cramér-Rao Bound
(4) $B=\theta-E\hat{\theta}_{X}=\int\big(\hat{\theta}_{X}-\theta\big)p_{\theta}(x)dx=0\quad\forall\theta,$
@@
(5) $E(\theta-\hat{\theta}_{X})^{2}\geq\frac{1}{J_{X}(\theta)},$
In other words, the Fisher information of the likelihood function fundamentally bounds the precision to which we can estimate $\theta$, for unbiased estimators.
@@ 2.5 Problem Formulation
Let $Y$ be an observation, and $\hat{\theta}_{Y}$ be the estimated parameter given $m$ examples (realizations) of $Y$. We assume $\hat{X}=\mathcal{F}(Y)$ is an estimated signal that is derived from $Y$ by solving an inverse problem, in an attempt to recover an underlying source signal $X$, and $\hat{\theta}_{\hat{X}}$ as the estimated parameter given $m$ examples of $\hat{X}$. In other words, $\theta-X-Y$ is a Markov chain. That is, given $X$, $Y$ and $\theta$ are independent. Namely, $p_{\theta}(x,y)\triangleq p(x,y|\theta)=p_{\theta}(x)p(y|x)$ (i.e., the conditional distribution of $Y$ given $X$ is independent of $\theta$).
@@
Definition 2.2.
A degradation $p_{Y|X}(\cdot|x)$ is invertible if the conditional probability distribution $P_{X|Y}(\cdot|y)$ is a delta measure for almost every $y$.
In other words, given the observation $Y=y$, the original signal $X$ is uniquely determined (almost surely). A degradation is invertible if there exists a measurable function $\mathcal{G}:\mathcal{Y}\rightarrow\mathcal{X}$ such that $Pr[X=\mathcal{G}(Y)]=1$.
@@
Figure 1: Illustration of an example of the problem setting: Classification with or without image-to-image regression. We observe a degraded image $y\sim p_{Y|\theta}$ according to some conditional probability density function $p_{Y|X}$, where the original image is a realization from a probability density function $x\sim p_{X|\theta}$.
@@ 3 Towards mathematical theory of domain variation
Throughout this section, we consider a discrete setting where $X\in\mathcal{X}^{n\times 1}$ and $Y\in\mathcal{Y}^{d\times 1}$ are discrete random variables.
Definition 3.1.
In a discrete setting, a degradation $p_{Y|X}(\cdot|x)$ is invertible if the conditional probability mass function $p_{X|Y}(\cdot|y)$ is a delta Kronecker for all $y\in\mathrm{supp}(p_{Y})$.
@@
The Domain-Shift Problem. In learning theory, domain shift is a variation of the data distribution between the source domain (training dataset) and target domain (the test dataset). Supervised learning models have demonstrated remarkable results in image-to-image translation tasks across numerous applications [Elad:2023, Pereg:2023B, Pereg:2024B]. However, it has been observed that while sufficient results were obtained for the training domain (source domain), these models collapse when tested on unseen target domains, such as under contrast and resolution variations [Pan:2021]. One of the most common techniques to improve generalization is data augmentation, normally accomplished by applying simple spatial and intensity transforms to the training data. Another popular strategy to improve generalization is domain randomization [Tobin:2017] attempting to significantly enlarge the training distribution beyond the source domain, under the hypothesis that learning from out-of-distribution examples improves generalization [Bengio:2011]. Most domain adaptation and transfer learning methods operate under the assumption that there exists a shared hypothesis that performs well in both domains [Ben:2010]. Here, we argue that certain implementations of this strategy may lead to degraded results, because instead of learning out-of-distribution examples (extrapolation), the model learns new examples within the distribution’s support that compete with the desired task and create ambiguity, which in turn leads to mode-collapse.
@@ 3.1 Signal Restoration over Several Domains
(6) $p_{\theta_{i}}(\mathbf{x}_{i}|\mathbf{y})=\delta(\mathbf{x}_{i}-G_{i}(\mathbf{y})),\ G_{i}(\mathbf{y})\neq G_{j}(\mathbf{y})\ \ \forall i\neq j,\ i,j=1,...,M,$
[…]
(7) $\arg\min_{\hat{\mathbf{x}}}\sum_{i=1}^{M}{\ell(\hat{\mathbf{x}},\mathbf{x}_{i})},$
@@
(8) $\hat{\mathbf{x}}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{x}_{i}.$
For the $\ell_{1}$ loss, the sum of absolute deviations $\ell(\hat{\mathbf{x}},\mathbf{x}_{i})=|\hat{\mathbf{x}}-\mathbf{x}_{i}|$, the optimum is at the median of the expected outputs. This phenomenon is also referred to in the literature as “regression to the mean” [Elad:2023], but normally in the context of having multiple restoration solutions in one domain due to the ill-posed nature of the inversion task (see also Section 4.3.2 below). Here, we emphasize that even if the restoration function is invertible in each domain  [^3] That is, in each domain the posterior is a delta function with nonzero probability for only one reconstruction. Therefore it is possible to perfectly reconstruct the signal from the measurement. unless the sets are disjoint, domain randomization would directly lead to averaging over the possible outputs. Let us summarize the above conclusion in the following theorem,
Theorem 3.2 (Double Meaning Theorem).
Let $p(\mathbf{x}|\theta_{i}),\ i=1,...,M$ be M domains, and $p(\mathbf{y}|{\mathbf{x}})$ be a degradation process, such that the observation sets are not disjoint, i.e., $p_{\theta_{i}}(\mathbf{x}|\mathbf{y})$ have joint supports in the $\mathbf{y}$ input domain. Let $\mathcal{F}(\mathbf{y})$ be an inversion model that is successfully trained by domain randomization. That is, the model is trained to minimize a cost function over all given and equiprobable domains, $P(\theta_{i})=\frac{1}{M},\ \forall i=1,...,M$, such that the model $\mathcal{F}$ is trained by minimizing the empirical loss
(9) $\mathcal{R}_{m}(\mathcal{F})=\frac{1}{m}\sum_{k=1}^{m}\ell(\mathcal{F}(\mathbf{y}_{k}),\mathbf{x}_{k}).$
where $m$ is the sample size. For $m$ large enough, the inversion output yields the minimal average expected loss over the possible outputs in the different domains,
(10) $\hat{\mathbf{x}}=\mathcal{F}(\mathbf{y})=\arg\min_{\hat{\mathbf{x}}}\frac{1}{M}\sum_{i=1}^{M}E_{\mathbf{x}_{i}\sim\mathbf{x}|\mathbf{y},\theta_{i}}\big[\ell(\hat{\mathbf{x}},\mathbf{x}_{i})\big],$
If $p(\mathbf{x}|\mathbf{y},\theta_{i})$ collapses to a Dirac delta at $\mathbf{x}_{i}$ for each domain $i$, then we have
(11) $\hat{\mathbf{x}}=\mathcal{F}(\mathbf{y})=\arg\min_{\hat{\mathbf{x}}}\frac{1}{M}\sum_{i=1}^{M}{\ell(\hat{\mathbf{x}},\mathbf{x}_{i})}.$
Proof. See Appendix B.
@@
Remark 2. Note that when the sets $\{\mathbf{y}\sim p_{\theta_{i}}(\mathbf{y})\}_{i=1}^{M}$ are disjoint, it is possible to train a model in a randomized manner over these distributions to generalize over the general $p(\mathbf{x}|\mathbf{y})$ without yielding averaged results. This may be the case for certain implementations of segmentation over different domains [Dey:2024], and for the MNIST dataset that can be modeled as a Gaussiam mixture model, under the assumption that the Gaussians are “far enough” from each other (in which case the supports are approximately disjoint).
@@
Remark 4. A similar proof outline applies also when a model is trained with noisy or perturbed training data (which can be treated as an adversarial attack [Sofer:2025]) and for other regression tasks. In those cases the model converges to the mean, which, depending on the task is either beneficial or misleading (see Section 3.4 below, and [Pereg:2026]).
Simple is beautiful. In previous works, it has also been observed that synthetic data training may improve generalization [Pereg:2020, Dey:2024, Pereg:2024B]. In those works, the synthetic data is of simpler visual form, almost degenerated. This may be explained by the opposing approach. That is, one may claim that using synthetic data of extremely simplified forms we are significantly limiting the learned distribution, and rather than expanding it, we are contracting the example set which the model has seen and is able to recognize. And so, by accurately learning simpler representations - the general result is improved. The model learns the basic pattern that may align with many distributions, inherently learning the overlapping support shared across a greater number of possible varying distributions. An alternative theoretical explanation relying on the information theoretic method of types is presented in [Pereg:2023A].
We now provide an illustration of the core idea in the Double Meaning Theorem (Theorem 3.2) in two relatively simple yet pragmatic examples.
@@ 3.2 Example 1 - Resolution Shift
Assume a source image and a degraded image in vector form, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{N\times 1}$, such that, $\{\mathbf{x},\mathbf{y}\sim p_{\theta_{i}}(\mathbf{x},\mathbf{y}):\mathbf{y}=\mathbf{H}_{i}\mathbf{x}+\mathbf{w},\ i=\{1,2\}\}$. $\mathbf{H}_{i}$ is the convolution matrix of a Gaussian filter with standard deviation $\sigma_{i}$, where $\sigma_{2}=2\sigma_{1}=2$, and $\mathbf{w}$ is additional white Gaussian noise (WGN) of level $\sigma_{n}=\sqrt{2}$. To validate our theoretical results, we compare trained one-shot supervised learning models for the task of image deblurring [Pereg:2023B, Pereg:2024B]. Recall that for a Gaussian PSF, the convolution of two Gaussian with mean $\mu_{1},\mu_{2}$ and variance $\sigma^{2}_{1}$, $\sigma^{2}_{2}$ is a Gaussian with mean $\mu=\mu_{1}+\mu_{2}$ and variance $\sigma^{2}=\sigma^{2}_{1}+\sigma^{2}_{2}$. Therefore, under this model, there exist $\mathbf{x}_{1}\neq\mathbf{x}_{2}$ such that $\mathbf{y}=\mathbf{H}_{1}\mathbf{x}_{1}=\mathbf{H}_{2}\mathbf{x}_{2}$. Let us denote $\mathbf{y}_{1}=\mathbf{H}_{1}\mathbf{x}_{1},\ \mathbf{y}_{2}=\mathbf{H}_{2}\mathbf{x}_{2}$, where we have omitted the noise term for that sake of simplicity. An identical degraded input image obeys $\mathbf{y}_{1}=\mathbf{y}_{2}$. Therefore,
(12) $\mathbf{x}_{1}*\mathbf{h}_{1}=\mathbf{x}_{2}*\mathbf{h}_{2}=\mathbf{x}_{2}*\mathbf{h}_{1}*\mathbf{h}_{1\rightarrow 2},$
where $\mathbf{h}_{1\rightarrow 2}$ is a Gaussian filter with $\sigma^{2}_{1\rightarrow 2}=\sigma_{2}^{2}-\sigma_{1}^{2}$. That is, $\mathbf{x}_{1}=\mathbf{x}_{2}*\mathbf{h}_{1\rightarrow 2}$. Consequently, according to Theorem 3.2, the outcome of a model that optimizes the MSE over both outcomes and thus averaging over both viable reconstructions is,
(13) $\hat{\mathbf{x}}=\frac{1}{2}\mathbf{x}_{1}+\frac{1}{2}\mathbf{x}_{2}=\frac{1}{2}\mathbf{x}_{2}+\frac{1}{2}\mathbf{x}_{2}*\mathbf{h}_{1\rightarrow 2}=\frac{1}{2}[\mathbf{I+H}_{1\rightarrow 2}]\mathbf{x}_{2}.$
@@
To illustrate the Double Meaning theorem above, we present a simple example using a one-shot training framework previously described in [Pereg:2023B, Pereg:2024B]. Figures 3-4(a)-(b) show the source image and the degraded image for $\sigma_{2}=2$, respectively. The inverted results and the residuals (absolute value) are depicted in figures 3-4(c)-(d) and figures 3(e)-(f), respectively. We compare the result of a model that was trained with images of both degradation types (mixed training) with a model that was trained solely for the Gaussian blur with a matching $\sigma_{2}=2$ (targeted training). As can be visually observed, in accordance with the Double Meaning Theorem, mixed training yields inferior results compared with targeted training where the training matches with the blurring kernel. As can be seen, as expected, the model cannot adhere to both required resolution levels and it “collapses to the mean”.
@@ 3.3 Example 2 - Sampling Shift
Similarly, we investigate the generalization for speckle-suppression in optical coherence tomography [Pereg:2023B], and compare the ability of the model to generalize over images of two lateral sampling rates $f^{2}_{\mathrm{s}}=2f^{1}_{\mathrm{s}}$. We test a model that is trained with retinal images of both sampling rates, and a single matching sampling rate. Figures 5-6 show the speckled images, despeckled images obtained by applying Non-Local-Means (NLM) filtering [Cuartas:2018] used as ground truth, the despeckled images produced by the trained models, and the residuals. Clearly, mixing sampling rates during training yields degraded results. Note the increased error around the blood vessel shadow projections (darker vertical area) in Fig. 5. The blood vessel shadow is almost smeared completely in the mixed training example. We observe that, as expected, the models’ prediction is more loyal to the measurement than to the NLM ground truth. Additionally, we can observe crossed-hatched artifacts as a result of inconsistencies in training sampling rates. For these examples the differences between the outputs may often seem subtle, because the averaged terms are visually similar and their average may also seem close to the ground truth.
@@
3.4 Noisy Training Data
Noisy Targets
Let $\Psi=\{\mathbf{y}_{k},\tilde{\mathbf{x}}_{k}\}_{k=1}^{m}$ be a training set with noisy outputs, where $\tilde{\mathbf{x}_{k}}\in\mathcal{X}^{n\times 1}:E[\tilde{\mathbf{x}}|\mathbf{x}]=\mathbf{x}$ are sampled from $p(\tilde{\mathbf{x}}|\mathbf{x})$ and paired with $\mathbf{y}_{k}\in\mathcal{Y}^{d\times 1}$ by a deterministic or stochastic mapping as ground truth. Let $\hat{\mathbf{x}}=\mathcal{F}(\mathbf{y})$ be an inversion model that is trained to minimize a cost function, such that the model $\mathcal{F}$ is trained by minimizing the empirical loss
(14) $\mathcal{R}_{m}(\mathcal{F})=\frac{1}{m}\sum_{k=1}^{m}\ell(\mathcal{F}(\mathbf{y}_{k}),\tilde{\mathbf{x}}_{k}).$
where $m$ is the sample size. For $m$ large enough,
[…]
For $m\rightarrow\infty$ and a convex loss
[…]
$=E_{\mathbf{y}}E_{\mathbf{x}|\mathbf{y}}\Big[E_{\tilde{\mathbf{x}}|\mathbf{y}}\big[\ell(\mathcal{F}(\mathbf{y}),\tilde{\mathbf{x}}|\mathbf{x})\big]\Big]\geq E_{\mathbf{x},\mathbf{y}}\big[\ell(\mathcal{F}(\mathbf{y}),\mathbf{x})\big],$
which follows from the law of total expectation and Jensen inequality.
Corollary 3.3.
In the special case of MSE loss $\ell(\hat{\mathbf{x}},\tilde{\mathbf{x}})=\|\hat{\mathbf{x}}-\tilde{\mathbf{x}}\|_{2}^{2}$, the minimizer is the conditional expectation $\mathcal{F}(\mathbf{y})=E[{\tilde{\mathbf{x}}|\mathbf{y}}]$ [Hastie:2009]. And since the noise is unbiased $E[\tilde{\mathbf{x}}|\mathbf{x}]=\mathbf{x}$, the minimizer is the clean posterior mean $\mathcal{F}(\mathbf{y})=\arg\min_{\hat{\mathbf{x}}}E_{\mathbf{x},\mathbf{y}}\big[\ell(\hat{\mathbf{x}},\mathbf{x})\big]=E[\mathbf{x}|\mathbf{y}]$. In other words, for MSE loss, unbiased target noise does not affect the learned model.
Proof 3.4.
Using the law of total expectation,
$E[\tilde{\mathbf{x}}|\mathbf{y}]=E[E[\tilde{\mathbf{x}}|\mathbf{x}]|\mathbf{y}]=E[\mathbf{x}|\mathbf{y}].$